OpenStudy (wade123):

Find the minimum value of the function f(x) = x^2+9x-16

Parth (parthkohli):

So many ways to do this. Do you know calculus yet?

Parth (parthkohli):

First assuming that you don't... In quadratic functions in the form \(f(x) = ax^2 + bx + c\) and \(a>0\), the minimum value of \(f(x)\) occurs when \(x= -\frac{b}{2a}\). In other words, \(f\left(-\frac{b}{2a}\right)\) is the minimum value.

OpenStudy (abb0t):

You can either find the root using algebra and factor, or take the first derivative.

OpenStudy (wade123):

sorry! i lost internet connection but i solved it and got –2.500 @ParthKohli

OpenStudy (wade123):

negative 2.500, am i right?

Parth (parthkohli):

Hmm, seems not...

OpenStudy (wade123):

really? are you sure? @ParthKohli

OpenStudy (wade123):

oh wait i thought this was for a different problem, for this i got–36.250

Parth (parthkohli):

That's right.

OpenStudy (wade123):

negative 36.250

OpenStudy (wade123):

yes! thanks !

Parth (parthkohli):

No problem!